# An AI-powered Bayesian generative modeling approach for causal inference in observational studies

### v2

@@ Abstract
Causal inference in observational studies with high-dimensional covariates presents significant challenges. We introduce CausalBGM, an AI-powered Bayesian generative modeling approach that captures the causal relationship among covariates, treatment, and outcome variables. The core innovation of CausalBGM lies in its ability to estimate the individual treatment effect (ITE) by learning individual-specific distributions of a low-dimensional latent feature set (e.g., latent confounders) that drives changes in both treatment and outcome. This approach not only effectively mitigates confounding effects but also provides comprehensive uncertainty quantification, offering reliable and interpretable causal effect estimates at the individual level. CausalBGM adopts a Bayesian model and uses a novel iterative algorithm to update the model parameters and the posterior distribution of latent features until convergence. This framework leverages the power of AI to capture complex dependencies among variables while adhering to the Bayesian principles. Extensive experiments demonstrate that CausalBGM consistently outperforms state-of-the-art methods, particularly in scenarios with high-dimensional covariates and large-scale datasets. Its Bayesian foundation ensures statistical rigor, providing robust and well-calibrated posterior intervals. By addressing key limitations of existing methods, CausalBGM emerges as a robust and promising framework for advancing causal inference in modern applications in fields such as genomics, healthcare, and social sciences. CausalBGM project is maintained at the website https://causalbgm.readthedocs.io/.
@@ 1 Introduction
One central goal for causal inference in observational studies is to estimate the causal effect of one variable (e.g., treatment) on another (e.g., outcome) while accounting for covariates that represent all other measured variables (Rothman and Greenland, 2005; Pearl, 2009; Imbens and Rubin, 2015; Ding, 2024). Covariates are often high-dimensional for modern applications in genomics, economics, and healthcare (Prosperi et al., 2020; Davey Smith et al., 2020; Forastiere et al., 2021), which makes the covariate adjustment difficult due to the “curse of dimensionality” (D’Amour et al., 2021). Additionally, large sample sizes, as is often the case in those scenarios, can further complicate the process by making traditional methods computationally intensive and slow to converge, highlighting the need for developing scalable and effective causal inference method.
@@
To handle the issue of high-dimensional covariates, several dimension reduction methods have been proposed to alleviate the difficulty. For example, one of the most popular approaches is to do adjustment or matching based on the propensity score (Rubin, 1974; Rosenbaum and Rubin, 1983; Hirano and Imbens, 2004), which is a one-dimensional feature (e.g., a scalar), denoting the probability of receiving a particular treatment given observed covariates. These methods require fitting a propensity score model first, which is typically done by fitting a logistic regression or a machine learning model (Lee et al., 2010). Another type of dimension reduction method is sufficient dimension reduction (SDR) (Li, 1991, 1992), which projects covariates into a lower-dimensional space, assuming conditional independence of treatment and outcome given the projected features (Ghosh et al., 2021; Luo et al., 2017). However, SDR-based causal inference methods often restrict dimension reduction to be linear transformations and apply separate projections for each treatment value, limiting its applicability in settings with continuous treatments or complex dependencies. The latent factor approach has also been used as surrogate confounders to adjust for biases in causal effect estimation caused by unobserved confounders (Yuan and Qu, 2024).
@@
Recently, the rapid development of AI-powered causal inference approaches has shown promising results for causal effect estimation (Berrevoets et al., 2023; Lagemann et al., 2023). These AI-based approaches typically leverage deep learning techniques and demonstrate superior power in modeling complex dependency and estimation accuracy when the sample size is large. In particular, the Causal Encoding Generative Modeling approach, CausalEGM (Liu et al., 2024), developed by our group, combines anto-encoding and generative modeling to enable nonlinear, structured dimension reduction in causal inference. CausalEGM stands at the intersection of AI and causal inference and has been shown to provide superior performance for developing deep learning-based estimates for the structural equation modeling that describes the causal relations among variables.
@@
To address the above issues, we introduce CausalBGM, a AI-powered Bayesian Generative Modeling (BGM) framework for estimating causal effects in the presence of high-dimensional covariates. Compared to CausalEGM, the new CausalBGM removes the encoder function entirely and employs a fully Bayesian procedure to infer latent features (Figure 1B without the dotted arrow). By eliminating the encoder-decoder loop, CausalBGM guarantees a clear DAG structure that is consistent with statistical modeling principles. Both the latent variables and model parameters are drawn from probabilistic distributions rather than being deterministically encoded, allowing for the incorporation of prior information and the generation of posterior distributions that more accurately represent uncertainty. By leveraging this fully Bayesian methodology, CausalBGM achieves substantial improvements, providing a principled alignment with Bayesian causal inference (Li et al., 2023). The model eliminates problematic cycles, adopts Bayesian inference, and ultimately provides a more robust and interpretable framework for estimating causal effects in complex, high-dimensional data settings. We highlight several key innovations of CausalBGM as follows.
@@ 2.2 Causal Generative Modeling
Note that the learnable functions ($\mu_{x},\sigma^{2}_{x}$),($\mu_{y},\sigma^{2}_{y}$), and ($\mu_{v},\sigma^{2}_{v}$) are represented by three Bayesian neural networks (Jospin et al., 2022), parameterized by $\theta_{X}$, $\theta_{Y}$, and $\theta_{V}$ respectively. In the next section, we will illustrate how we learn the distribution of model parameters $\theta_{X}$, $\theta_{Y}$, $\theta_{V}$ in order to account for the model uncertainty or variation.
@@ 2.3 Iterative Updating Algorithm
where $\epsilon_{X}\sim\mathcal{N}(0,I_{d_{X}})$, $\epsilon_{Y}\sim\mathcal{N}(0,I_{d_{Y}})$, and $\epsilon_{V}\sim\mathcal{N}(0,I_{d_{V}})$. $\odot$ is the element-wise product. $d_{X}$, $d_{Y}$, and $d_{V}$ denote the number of parameters in each generative model. Using variational inference in BNNs for each mini-batch may lead to high-variance gradient estimates. We adopt the Flipout technique (Wen et al., 2018) in the implementation of the reparameterization trick to reduce this variance by decorrelating the model parameters perturbations across different training examples in the same mini-batch. Briefly, in stead of using a single shared random draw of model parameters for the entire mini-batch, Flipout constructs pseudo-independent perturbations for each example independently within a mini-batch, which decorrelates the gradients, reduces the variance, and stabilizes the training process.
@@
Each iteration only requires the current sample or a random mini-batch of observed samples. During the iteration algorithm, we first take a derivative of equation (10) w.r.t the latent variable $Z$ and employ a stochastic gradient descent (SGD) to update the latent variable $Z$ for each individual given the current model parameters. Then, we take a derivative of each ELBO term in (13) w.r.t the variational parameters ($\phi_{X}$, $\phi_{Y}$, or $\phi_{V}$) in the three generative models sequentially and employ a stochastic gradient ascent to update the variational parameters given the current latent variables to maximize the ELBO. During test stage, we only need to infer the posterior distribution of latent variable $Z$ given the test data. To achieve this, we first sampled model parameters ($\theta_{X}$, $\theta_{Y}$, and $\theta_{V}$) from the variational distribution parameterized by $\phi_{X}$, $\phi_{Y}$, and $\phi_{V}$ through (14). Then we use Markov chain Monte Carlo (MCMC) method (Liu, 2001) to sample from the posterior distribution in (10) for each individual. We choose the standard Metropolis–Hastings algorithm (Robert et al., 2004) as default. Note that this individual-level sampling process is fully decoupled, enabling parallelization and improving computational efficiency. The causal effect and the corresponding posterior interval with user-specified significant level, can be then estimated based on the MCMC samples of latent variable and the learned generative models.
@@ 2.4 Choice of latent dimension
The latent space is partitioned into four independent parts that play different roles in the three generative models for treatment, covariates, and outcome variables. The previous work CausalEGM has demonstrated the robustness of such partition with respect to variations in the dimensionality of latent features. Here, we provide an intuitive strategy based on sufficient dimension reduction (SDR) to help determine the dimensionality of latent features. SDR aims to identify a $k$-dimensional subspace of the $p$-dimensional predictors ($k\ll p$) that captures all the information about a scalar response. Here, we use sliced inverse regression (SIR) (Li, 1991) that employs the covariance structure of the conditional expectations of predictors given response. We compute eigenvalues of the estimated covariance matrix from SIR and retain components by inspecting eigenvalue decay and cumulative variance. Considering that linear methods such as SIR may underestimate $k$ as they fail to capture nonlinear dependencies effectively. Here, we use a conservative strategy by using SIR $\mathbb{E}[V|X]$ to estimate $dim(Z_{2})$ and using SIR $\mathbb{E}[V|Y]$ to estimate $dim(Z_{1})$. A similar eigenvalue analysis for the covariance matrix of $V$ is conducted to estimate the total dimension of the latent space $dim(Z)$. The dimension of latent confounder $dim(Z_{0})$ is chosen from 1 to 5 as a model hyperparameter.
@@ 2.5 Model Initialization
The parameters of neural networks (e.g., weights and biases) are typically initialized by a uniform or normal distribution. However, as shown by our experiments (Table 3), the model performance can be further improved in most cases through our designed innovative strategy for model parameters initialization, inspired from the encoding generative modeling (EGM) (Liu et al., 2024), compared to the traditional neural network initialization. An additional encoder function $E$, represented by a Bayesian neural network, is added to CausalBGM to directly map the covariate $V$ to the latent variable (dotted line in Figure 1). Specifically, we desire that the distribution of $Z=E(V)$ should match a pre-specified distribution, which is set to be a standard normal distribution (e.g., prior of $Z$). The distribution match is achieved by adversarial training (Goodfellow et al., 2014). By the encoding process, the high-dimensional covariates with unknown distribution are mapped to a low-dimensional latent space with a desired distribution. Since the generative models in CausalBGM include both the mean and variance functions. We add additional The $L_{2}$ regularization of all the variance terms to ensure reasonably small variance in each generative model during the initialization process.
@@ 2.6 Model Hyperparameters
CausalBGM contains three generative models, which are represented by three Bayesian neural networks (BNNs), respectively. The BNN for covariate $V$ contains 5 hidden Bayesian layers and each layer has 64 hidden nodes. The output of BNN for covariate $V$ is $(p+1)$-dimensional where the first $p$ digits denote the mean $\mu_{v}$ and the last digit denotes variance $\sigma^{2}_{v}$. The BNN for treatment $X$ and outcome $Y$ contains 3 hidden Bayesian layers with 64, 32, and 8 hidden nodes. The output of BNN of treatment $X$ and outcome $Y$ is 2 dimensional, representing the mean and variance, respectively. The leaky-ReLu function ($LeakyReLU(x)=max(0.2x,x)$) is used as the non-linear activation function in each hidden layer. The Softplus non-linear activation function ($Softplus(x)=log(1+e^{x})$) is applied to the last digit of the BNN output to ensure positivity of variance. Adam optimizers (Kingma, 2014) with learning rate $0.0001$ are used to update latent variable and model parameters, respectively. The CausalBGM model is trained in a mini-batch manner with batch size 32. The default training epochs of CausalBGM with random initilization strategy is 500. If EGM initilization strategy is used, we initialize model parameters of CausalBGM by conducting EGM for $30,000$ mini-batches as default. After model initialization, the encoder $E$ as a “shortcut” to learn the latent variable is removed during the follow-up CausalBGM training with an iterative approach for up to 100 epochs. In the random walk Metropolis algorithm, we set the proposal distribution to be a normal distribution centered at the current sample with covariance matrix $I_{q}$. The Markov chain samples from the first $5,000$ iterations are discarded so that the effect of initial values is minimized (burn-in stage). Then we run the Markov chains parallelly for all samples until $3,000$ MCMC latent samples are collected for each sample.
@@ 3.1 Datasets
For the continuous treatment setting, we examined four public datasets used in previous studies (Hirano and Imbens, 2004; Sun et al., 2015; Colangelo and Lee, 2020), comprising three simulated datasets and one semi-synthetic dataset. Each of the simulation datasets has $20,000$ as the sample size and $200$ covariate features. We focus on the ADRF estimate in a bounded interval. The semi-synthetic data were derived from a sample of 71,345 twin births, where weight served as the continuous treatment variable and the risk of death is treated as the outcome variable, which is simulated from a risk model. Each individual has 50 covariates. In general, the simulation risk model suggests that a higher weight of an infant leads to a lower death rate.
@@ 3.3 Baseline Methods
For the continuous treatment setting, we considered four well-established baseline methods: ordinary least squares (OLS), the regression prediction estimator (REG) (Schafer and Galagate, 2015; Galagate, 2016; Imai and Van Dyk, 2004), double debiased machine learning estimators (DML) (Colangelo and Lee, 2020), and CausalEGM (Liu et al., 2024). Note that different machine learning methods shall be used in the DML method. For the binary treatment setting, we compared CausalBGM against seven leading methods for estimating treatment effect, including two variants of CFR (Shalit et al., 2017), Dragonnet (Shi et al., 2019), CEVAE (Louizos et al., 2017), GANITE (Yoon et al., 2018), CausalForest (Wager and Athey, 2018), and CausalEGM (Liu et al., 2024). Additional details about these competing methods are provided in Appendix D.

### v3

@@ Abstract
Causal inference in observational studies with high-dimensional covariates presents significant challenges. We introduce CausalBGM, an AI-powered Bayesian generative modeling approach that captures the causal relationship among covariates, treatment, and outcome. The core innovation is to estimate the individual treatment effect (ITE) by learning the individual-specific distribution of a low-dimensional latent feature set (e.g., latent confounders) that drives changes in both treatment and outcome. This individualized posterior representation yields estimates of the individual treatment effect (ITE) together with well-calibrated posterior intervals while mitigating confounding effect. CausalBGM is fitted through an iterative algorithm to update the model parameters and the latent features until convergence. This framework leverages the power of AI to capture complex dependencies among variables while adhering to the Bayesian principles. Extensive experiments demonstrate that CausalBGM consistently outperforms state-of-the-art methods, particularly in scenarios with high-dimensional covariates and large-scale datasets. By addressing key limitations of existing methods, CausalBGM emerges as a robust and promising framework for advancing causal inference in a wide range of modern applications. The code for CausalBGM is available at https://github.com/liuq-lab/bayesgm. The tutorial for CausalBGM is available at https://causalbgm.readthedocs.io. Supplementary materials for this article are available online.
@@ 1 Introduction
One central goal for causal inference in observational studies is to estimate the causal effect of one variable (e.g., treatment) on another (e.g., outcome) while accounting for covariates that represent all other measured variables (rothman2005causation; Pearl2009-PEACII; imbens2015causal; ding2024first). Covariates are often high-dimensional for modern applications in genomics, economics, and healthcare (prosperi2020causal; davey2020mendel; forastiere2021identification), which makes the covariate adjustment difficult due to the “curse of dimensionality” (d2021overlap). Additionally, large sample sizes, as is often the case in those scenarios, can further complicate the process by making traditional methods computationally intensive and slow to converge, highlighting the need for developing scalable and effective causal inference method.
@@
To handle the issue of high-dimensional covariates, several dimension reduction methods have been proposed to alleviate the difficulty. For example, one of the most popular approaches is to do adjustment or matching based on the propensity score (rubin1974estimating; rosenbaum_rubin1983; hirano2004propensity), which is a one-dimensional feature (e.g., a scalar), denoting the probability of receiving a particular treatment given observed covariates. These methods require fitting a propensity score model first, which is typically done by fitting a logistic regression or a machine learning model (lee2010improving). Another type of dimension reduction method is sufficient dimension reduction (SDR) (li1991sliced; li1992principal), which projects covariates into a lower-dimensional space, assuming conditional independence of treatment and outcome given the projected features (ghosh2021sufficient; luo2017estimating). However, SDR-based causal inference methods often restrict dimension reduction to be linear transformations and apply separate projections for each treatment value, limiting its applicability in settings with continuous treatments or complex dependencies. The latent factor approach has also been used as surrogate confounders to adjust for biases in causal effect estimation caused by unobserved confounders (yuan2024confounding).
@@
Recently, the rapid development of AI-powered causal inference approaches has shown promising results for causal effect estimation (berrevoets2023causal; lagemann2023deep). These AI-based approaches typically leverage deep learning techniques and demonstrate superior power in modeling complex dependency and estimation accuracy when the sample size is large. In particular, the Causal Encoding Generative Modeling approach, CausalEGM (Liu_pnas2024), developed by our group, combines anto-encoding and generative modeling to enable nonlinear, structured dimension reduction in causal inference. CausalEGM stands at the intersection of AI and causal inference and has been shown to provide superior performance for developing deep learning-based estimates for the structural equation modeling that describes the causal relations among variables.
@@
To address the above issues, we introduce CausalBGM, a AI-powered Bayesian Generative Modeling (BGM) framework for estimating causal effects in the presence of high-dimensional covariates. Compared to CausalEGM, the new CausalBGM removes the encoder function entirely and employs a fully Bayesian procedure to infer latent features (Figure 1B without the dotted arrow). By eliminating the encoder-decoder loop, CausalBGM guarantees a clear DAG structure that is consistent with statistical modeling principles. Both the latent variables and model parameters are drawn from probabilistic distributions rather than being deterministically encoded, allowing for the incorporation of prior information and the generation of posterior distributions that more accurately represent uncertainty. By leveraging this fully Bayesian methodology, CausalBGM achieves substantial improvements, providing a principled alignment with Bayesian causal inference (li2023bayesian). The model eliminates problematic cycles, adopts Bayesian inference, and ultimately provides a more robust and interpretable framework for estimating causal effects in complex, high-dimensional data settings. We highlight several key innovations of CausalBGM as follows.
@@ 2.2 Causal Generative Modeling
Note that the learnable functions ($\mu_{x},\sigma^{2}_{x}$),($\mu_{y},\sigma^{2}_{y}$), and ($\mu_{v},\sigma^{2}_{v}$) are represented by three Bayesian neural networks (jospin2022hands), parameterized by $\theta_{X}$, $\theta_{Y}$, and $\theta_{V}$ respectively. In the next section, we will illustrate how we learn the distribution of model parameters $\theta_{X}$, $\theta_{Y}$, $\theta_{V}$ in order to account for the model uncertainty or variation.
@@ 2.3 Iterative Updating Algorithm
where $\epsilon_{X}\sim\mathcal{N}(0,I_{d_{X}})$, $\epsilon_{Y}\sim\mathcal{N}(0,I_{d_{Y}})$, and $\epsilon_{V}\sim\mathcal{N}(0,I_{d_{V}})$. $\odot$ is the element-wise product. $d_{X}$, $d_{Y}$, and $d_{V}$ denote the number of parameters in each generative model. Using variational inference in BNNs for each mini-batch may lead to high-variance gradient estimates. We adopt the Flipout technique (wen2018flipout) in the implementation of the reparameterization trick to reduce this variance by decorrelating the model parameters perturbations across different training examples in the same mini-batch. Briefly, in stead of using a single shared random draw of model parameters for the entire mini-batch, Flipout constructs pseudo-independent perturbations for each example independently within a mini-batch, which decorrelates the gradients, reduces the variance, and stabilizes the training process.
@@
Each iteration only requires the current sample or a random mini-batch of observed samples. During the iteration algorithm, we first take a derivative of equation (10) w.r.t the latent variable $Z$ and employ a stochastic gradient descent (SGD) to update the latent variable $Z$ for each individual given the current model parameters. Then, we take a derivative of each ELBO term in (13) w.r.t the variational parameters ($\phi_{X}$, $\phi_{Y}$, or $\phi_{V}$) in the three generative models sequentially and employ a stochastic gradient ascent to update the variational parameters given the current latent variables to maximize the ELBO. During test stage, we only need to infer the posterior distribution of latent variable $Z$ given the test data. To achieve this, we first sampled model parameters ($\theta_{X}$, $\theta_{Y}$, and $\theta_{V}$) from the variational distribution parameterized by $\phi_{X}$, $\phi_{Y}$, and $\phi_{V}$ through (14). Then we use Markov chain Monte Carlo (MCMC) method (liu2001monte) to sample from the posterior distribution in (10) for each individual. We choose the standard Metropolis–Hastings algorithm (robert2004metropolis) as default. Note that this individual-level sampling process is fully decoupled, enabling parallelization and improving computational efficiency. The causal effect and the corresponding posterior interval with user-specified significant level, can be then estimated based on the MCMC samples of latent variable and the learned generative models.
@@ 2.4 Choice of latent dimension
The latent space is partitioned into four independent parts that play different roles in the three generative models for treatment, covariates, and outcome variables. The previous work CausalEGM has demonstrated the robustness of such partition with respect to variations in the dimensionality of latent features. Here, we provide an intuitive strategy based on sufficient dimension reduction (SDR) to help determine the dimensionality of latent features. SDR aims to identify a $k$-dimensional subspace of the $p$-dimensional predictors ($k\ll p$) that captures all the information about a scalar response. Here, we use sliced inverse regression (SIR) (li1991sliced) that employs the covariance structure of the conditional expectations of predictors given response. We compute eigenvalues of the estimated covariance matrix from SIR and retain components by inspecting eigenvalue decay and cumulative variance. Considering that linear methods such as SIR may underestimate $k$ as they fail to capture nonlinear dependencies effectively. Here, we use a conservative strategy by using SIR $\mathbb{E}[V|X]$ to estimate $dim(Z_{2})$ and using SIR $\mathbb{E}[V|Y]$ to estimate $dim(Z_{1})$. A similar eigenvalue analysis for the covariance matrix of $V$ is conducted to estimate the total dimension of the latent space $dim(Z)$. The dimension of latent confounder $dim(Z_{0})$ is chosen from 1 to 5 as a model hyperparameter.
@@ 2.5 Model Initialization
The parameters of neural networks (e.g., weights and biases) are typically initialized by a uniform or normal distribution. However, as shown by our experiments (Table 3), the model performance can be further improved in most cases through our designed innovative strategy for model parameters initialization, inspired from the encoding generative modeling (EGM) (Liu_pnas2024), compared to the traditional neural network initialization. An additional encoder function $E$, represented by a Bayesian neural network, is added to CausalBGM to directly map the covariate $V$ to the latent variable (dotted line in Figure 1). Specifically, we desire that the distribution of $Z=E(V)$ should match a pre-specified distribution, which is set to be a standard normal distribution (e.g., prior of $Z$). The distribution match is achieved by adversarial training (goodfellow2014generative). By the encoding process, the high-dimensional covariates with unknown distribution are mapped to a low-dimensional latent space with a desired distribution. Since the generative models in CausalBGM include both the mean and variance functions. We add additional The $L_{2}$ regularization of all the variance terms to ensure reasonably small variance in each generative model during the initialization process.
@@ 2.6 Model Hyperparameters
CausalBGM contains three generative models, which are represented by three Bayesian neural networks (BNNs), respectively. The BNN for covariate $V$ contains 5 hidden Bayesian layers and each layer has 64 hidden nodes. The output of BNN for covariate $V$ is $(p+1)$-dimensional where the first $p$ digits denote the mean $\mu_{v}$ and the last digit denotes variance $\sigma^{2}_{v}$. The BNN for treatment $X$ and outcome $Y$ contains 3 hidden Bayesian layers with 64, 32, and 8 hidden nodes. The output of BNN of treatment $X$ and outcome $Y$ is 2 dimensional, representing the mean and variance, respectively. The leaky-ReLu function ($LeakyReLU(x)=max(0.2x,x)$) is used as the non-linear activation function in each hidden layer. The Softplus non-linear activation function ($Softplus(x)=log(1+e^{x})$) is applied to the last digit of the BNN output to ensure positivity of variance. Adam optimizers (kingma2014adam) with learning rate $0.0001$ are used to update latent variable and model parameters, respectively. The CausalBGM model is trained in a mini-batch manner with batch size 32. The default training epochs of CausalBGM with random initilization strategy is 500. If EGM initilization strategy is used, we initialize model parameters of CausalBGM by conducting EGM for $30,000$ mini-batches as default. After model initialization, the encoder $E$ as a “shortcut” to learn the latent variable is removed during the follow-up CausalBGM training with an iterative approach for up to 100 epochs. In the random walk Metropolis algorithm, we set the proposal distribution to be a normal distribution centered at the current sample with covariance matrix $I_{q}$. The Markov chain samples from the first $5,000$ iterations are discarded so that the effect of initial values is minimized (burn-in stage). Then we run the Markov chains parallelly for all samples until $3,000$ MCMC latent samples are collected for each sample.
@@ 3.1 Datasets
For the continuous treatment setting, we examined four public datasets used in previous studies (hirano2004propensity; sun2015causal; colangelo2020double), comprising three simulated datasets and one semi-synthetic dataset. Each of the simulation datasets has $20,000$ as the sample size and $200$ covariate features. We focus on the ADRF estimate in a bounded interval. The semi-synthetic data were derived from a sample of 71,345 twin births, where weight served as the continuous treatment variable and the risk of death is treated as the outcome variable, which is simulated from a risk model. Each individual has 50 covariates. In general, the simulation risk model suggests that a higher weight of an infant leads to a lower death rate.
@@ 3.3 Baseline Methods
For the continuous treatment setting, we considered four well-established baseline methods: ordinary least squares (OLS), the regression prediction estimator (REG) (schafer2015causal; galagate2016causal; imai2004causal), double debiased machine learning estimators (DML) (colangelo2020double), and CausalEGM (Liu_pnas2024). Note that different machine learning methods shall be used in the DML method. For the binary treatment setting, we compared CausalBGM against seven leading methods for estimating treatment effect, including two variants of CFR (shalit2017estimating), Dragonnet (shi2019adapting), CEVAE (louizos2017causal), GANITE (yoon2018ganite), CausalForest (wager2018estimation), and CausalEGM (Liu_pnas2024). Additional details about these competing methods are provided in Appendix D.